OpenStudy (anonymous):

Solve for x, where x is a real number. Show work: x^2+3x-4=14

OpenStudy (anonymous):

x^2 +3x -18=0 (x +6)(x-3)=0 x=-6,3

jimthompson5910 (jim_thompson5910):

Solutions are x=3 or x=-6 See attached for work on how I got those solutions.

OpenStudy (anonymous):

sweet thanks. i can work through the other ones with an example. thanks!